OpenStudy (anonymous):

solve the system by substitution... show work -2x-y=-14 3x-y=11

OpenStudy (anonymous):

x==0, y==-11

OpenStudy (anonymous):

how did you get that?

OpenStudy (dumbcow):

solve one equation for y: 3x-y =11 --> y = 3x -11 substitute it into other equation -replace y with (3x-11) -2x-y = -14 --> -2x -(3x-11) = -14 solve for x -2x -3x +11 = -14 -5x = -25 x = 5 Substitute 5 in for x to solve for y: y=3x-11 --> y = 3(5) -11 y = 15 -11 y = 4